# CuClarabel: GPU Acceleration for a Conic Optimization Solver

### v2

@@ 4.3 Second-order cone programming
We compare our GPU implementation with the CPU-based Clarabel solver and Mosek solver. We also include results of the Mosek solver with pre-solve for comparison, which is denoted as Mosek* in the plots. The maximum termination time is again set to $1$h. We benchmark the second-order cone relaxations of 120 problems from the PGLib-OPF library, where the number of second-order cones exceed $2000$.
@@
Overall, the GPU solver is several times faster than Rust-based CPU solver. However, note that ClarabelGPU fails on two examples that ClarabelRs successfully solves. This highlights the different numerical performance of the linear system solvers between CUDSS and faer.
@@
Shifted GM Full Acc. 1.0 6.63 3.62 34.33
Low Acc. 1.0 7.02 2.99 36.36
Failure Rate (%) Full Acc. 1.7 39.0 13.6 90.7
Low Acc. 0.0 39.0 3.4 90.7
@@ 4.4 Exponential cone programming
We benchmark two classes of exponential cone programming problems with varying dimensionality, the logistic regression and the entropy maximization problems [13].
Logistic regression is a classical machine learning model for binary classification. Given a feature matrix $A\in\textbf{R}^{m\times n}$ and a vector of labels $b\in\textbf{R}^{m}$ such that $b_{i}\in\{0,1\}$, we formulate the sparse logistic regression problem as:
$\text{minimize}\ \ \sum_{i=1}^{m}\left(\log\left(1+\exp\left(x^{T}a_{i}\right)\right)-b_{i}x^{T}a_{i}\right)+\lambda\|x\|_{1},$
where $x$ is our variable and $a_{i}^{T}$ is the $i$-th row of $A$. Note that this is the sparse logistic regression problem due to the $\ell_{1}$-norm penalty on $x$. Indeed, when $\lambda=0$, we recover the classical logistic regression problem.
Each entry of $A$ is generated via $A_{ij}\sim\mathcal{N}(0,1)$. The label $b$ is generated following the same method in [57, 13]. We set $m=5n$ and vary value of $n$ from $2000$ to $10000$.
The entropy maximization problem aims to maximize entropy over a probability distribution given a set of $m$ linear inequality constraints, which can be interpreted as bounds on the expectations of arbitrary functions. The problem is formulated as
@@
Each element of $A$ is generated from the distribution $A_{ij}\sim\mathcal{N}(0,n)$. Then, we set $b=Av/1^{\top}v$ where $v\in\textbf{R}^{n}$ is generated randomly from $v_{i}\sim U[0,1]$, which ensures the problem is always feasible. We set $m$ to the nearest integer of $0.5n$ and vary value of $n$ from $2000$ to $10000$.
The benchmark results for these two problems with varying dimensionality are shown in Figure 4. From §D, even though the GPU acceleration of Clarabel is offset by nearly doubled number of iterations compared to Mosek, we can still achieve more than 2 times of acceleration for the overall time. That is to say we can possibly achieve more acceleration in ClarabelGPU if we can improve the numerical stability to the same level of Mosek on exponential cone programs. Compared to ClarabelRs, we find ClarabelGPU can benefit from GPU computation up to 20x times faster on logistic regression problems.
Figure 4: Performance profiles for the exponential cone programming problem set
(a) Relative performance profile
(b) Absolute performance profile
ClarabelGPU Mosek* ClarabelRs Mosek
Shifted GM Full Acc. 1.0 1.47 8.42 1.58
Low Acc. 1.0 1.47 8.42 1.58
Failure Rate (%) Full Acc. 0.0 0.0 40.0 0.0
Low Acc. 0.0 0.0 20.0 0.0
(c) Benchmark timings as shifted geometric mean and failure rates
@@ 4.5 Mixed precision
We test the mixed precision setting on QPs, including portfolio optimization and Huber fitting problems from §4.2. We compare it with the standard GPU implementation of Float64 data type. The results in Table LABEL:table:mixed_large_qp show that the use of mixed precision strategy can reduce the solve time up to a factor of $2$. This will be beneficial to problems like the parametric programming, where the setup is needed only once and the symbolic factorization structure can be reused repeatedly later on [58]. However, we also notice that the mixed precision may not be sufficient for conic optimization problems where the Hessian block $H$ in (7) is ill-conditioned, especially for the exponential cone that the exponential component causes huge variation in the magnitude of different entries.
@@
iterations total time (s)
@@
portfolio_optimization_n_10000 19 19 0.729 0.631
portfolio_optimization_n_25000 19 21 7.08 4
portfolio_optimization_n_20000 19 23 3.91 2.54
huber_fitting_n_5000 9 9 2.61 1.32
huber_fitting_n_25000 10 10 143 58.3
huber_fitting_n_10000 10 10 12.6 5.76
portfolio_optimization_n_5000 17 18 0.196 0.259
huber_fitting_n_15000 10 10 34.9 15.1
portfolio_optimization_n_15000 19 19 1.88 1.18
huber_fitting_n_20000 9 9 69.8 29.1
@@ 5 Conclusion
We have developed a GPU interior point solver for conic optimization. [^2] https://github.com/cvxgrp/CuClarabel We propose a mixed parallel computing strategy to process linear constraints with second-order cone, exponential cone and power cone constraints. The GPU solver shows several times of acceleration compared to state-of-the-art CPU conic solvers on many problems to high precision, including QPs, SOCPs and exponential programmings.
We also note that some initialization steps, such as population of the KKT matrix and matrix equilibration, take non-negligible time for large problems. We will also move the setup steps onto a GPU in our future development. The proposed mixed parallel computing strategy for GPU implementation is also applicable for conic solvers based on first-order operator-splitting methods.
@@ Acknowledgments
The work of Yuwen Chen was funded by EPSRC Impact Acceleration Account Fund. Parth Nobel was supported in part by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1656518. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The work of Stephen Boyd was supported by ACCESS (AI Chip Center for Emerging Smart Systems), sponsored by InnoHK funding, Hong Kong SAR.
Appendix A Supported cones in Clarabel and CuClarabel
We support the following atomic cones in our GPU solver:
  • The zero cone, defined as
    The dual cone of the zero cone is ($\left(\{0\}^{n}\right)^{*}=\textbf{R}^{n}$.
  • The nonnegative cone, defined as
  • The exponential cone, a $3$-dimensional cone defined as
    with its dual cone given by
    with its dual cone given by
Appendix B Barrier functions for different cones
For cones introduced in §A, the corresponding logarithmically-homogeneous self-concordant barrier functions with degree $\nu$ are given by:
  1. 1.
    Nonnegative cone $\textbf{R}_{+}^{n}$ of degree $n$:
    $f(z)=-\sum_{i\in\llbracket n\rrbracket}\log(z_{i}),\ z\in\textbf{R}_{+}^{n}.$
  2. 2.
    Second order cone $\mathcal{K}_{q}^{n}$ of degree $1$:
    $f(z)=-\frac{1}{2}\log\left(z_{1}^{2}-\sum_{i=2}^{n}z_{i}^{2}\right),\ z\in\mathcal{K}_{q}^{n}.$
  3. 3.
    Dual exponential cone $\mathcal{K}_{\mathrm{exp}}^{*}$ of degree $3$:
    $f(z)=-\log\left(z_{2}-z_{1}-z_{1}\log\left(\frac{z_{3}}{-z_{1}}\right)\right)-\log(-z_{1})-\log(z_{3}),\ z\in\mathcal{K}_{\mathrm{exp}}^{*}.$
  4. 4.
    Dual power cone $\mathcal{K}_{\mathrm{pow}}^{*}$ of degree $3$:
    $f(z)=-\log\left(\left(\frac{z_{1}}{\alpha}\right)^{2\alpha}\left(\frac{z_{2}}{1-\alpha}\right)^{2(1-\alpha)}-z_{3}^{2}\right)-(1-\alpha)\log(z_{1})-\alpha\log(z_{2}),\ z\in\mathcal{K}_{\mathrm{pow}}^{*}.$
Appendix C Scaling matrices
@@
The matrix $H(w)$ can be factorized as $H^{-1}(w)=W^{\top}W$, and we set $H=H^{-1}(w)$ in (6). The factors $w,W$ are then computed following [47].
@@
Algorithm 1 shows how to update the scaling matrix $H$ at iteration $k+1$:
Algorithm 1 Update the scaling matrix $H$
1:  Input: the scaling matrix $H^{k}$ at iteration $k$
2:
3:  Update $H_{1}^{k+1}$   // Zero cone
4:  Update $H_{2}^{k+1}$   // Nonnegative cone
5:
6:  @sync for $t=3$ to $i$   // Second-order cones
7:  $H_{t}^{k+1}=$soc_update_H($H_{t}^{k}$)
8:  end
@@
10:  @sync for $t=i+1$ to $j$   // Exponential cones
11:  $H_{t}^{k+1}=$exp_update_H($H_{t}^{k}$)
12:  end
13:
14:  @sync for $t=j+1$ to $p$   // Power cones
15:  $H_{t}^{k+1}=$pow_update_H($H_{t}^{k}$)
16:  end
17:
18:  return  $H^{k+1}$   // Output the new scaling matrix $H^{k+1}$
Appendix D Detailed benchmark results
Table 2: Solve times and iteration counts for the large LPs problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU Mosek ClarabelGPU Mosek
qap15 28605 22275 117225 0 22 16 4.72 3.6
brazil3 62488 23968 181026 0 20 15 0.66 1.53
ex10 104968 17680 1197360 0 13 16 5.75 14.7
neos_5052403_cygnet 104004 32868 4964040 0 23 12 7.66 4.86
supportcase10 195224 14770 584622 0 35 20 9.6 5.65
square41 164591 62234 13690857 0 28 12 8.39 28.5
chromaticindex1024_7 215039 73728 417780 0 9 10 1.8 6.4
neos_3025225 231263 69846 9497642 0 28 19 50.4 19.9
rmine15 443271 42438 964608 0 55 29 41.2 14
physiciansched3_3 417922 79555 1214174 0 452 284 251 174
neos_5251015 486996 136971 1955853 0 - 9 - 22.6
graph40_40 566100 102600 1466100 0 11 - 3.36 -
datt256_lp 535365 262144 2028020 0 10 5 7.52 223
s250r10 557246 273142 1864891 0 103 32 29.3 6.56
s100 743567 364417 2506751 0 92 33 96.1 9.42
savsched1 877295 328575 2351813 0 28 16 10.9 85.8
woodlands09 958865 382147 3410269 0 - 8 - 344
scpm1 1005000 500000 7250000 0 27 24 24.4 13.2
bharat 2030687 590519 4296196 0 67 - 67.3 -
tpl_tub_ws1617 2649817 747601 6215769 0 141 63 110 133
fhnw_binschedule1 3056177 1141653 10894631 0 49 - 232 -
supportcase19 2868909 1429098 7145290 0 64 17 56 12.5
s82 3469140 1690631 10403870 0 - 71 - 235
Table 3: Solve times and iteration counts for the QPs problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU ClarabelRs Gurobi Mosek ClarabelGPU ClarabelRs Gurobi Mosek
portfolio_optimization_n_5000 5501 5500 1259971 5500 15 15 17 9 0.442 7.9 0.9 2.65
portfolio_optimization_n_10000 11001 11000 5022139 11000 18 18 19 7 1.67 57.1 4.53 10.7
portfolio_optimization_n_15000 16501 16500 11282198 16500 18 18 18 8 4.16 165 10.4 28.7
portfolio_optimization_n_20000 22001 22000 20040608 22000 17 17 18 8 8.06 372 22.3 63.1
huber_fitting_n_5000 22500 27500 4725756 7500 8 8 13 15 4.24 68.7 6.53 10.3
portfolio_optimization_n_25000 27501 27500 31300056 27500 18 18 20 9 14.7 600 44.2 124
huber_fitting_n_10000 45000 55000 18826645 15000 8 8 11 17 18.7 636 39.7 77
huber_fitting_n_15000 67500 82500 42304462 22500 8 8 11 18 48.6 2.18e+03 110 383
huber_fitting_n_20000 90000 110000 75138584 30000 8 - 14 20 99.2 - 269 1.21e+03
huber_fitting_n_25000 112500 137500 117388752 37500 8 - 11 19 180 - 400 2.31e+03
Table 4: Solve times and iteration counts for the large OPF LP problem set
iterations total time (s)
Problem vars. cons. nnz(A) nnz(P) ClarabelGPU ClarabelRs Mosek ClarabelGPU ClarabelRs Mosek
case10192_epigrids 92051 27914 180020 697 17 17 35 0.742 0.533 0.764
case10192_epigrids__api 92051 27914 180020 697 17 17 28 0.706 0.526 0.627
case10480_goc 99926 29816 196673 276 20 20 27 0.82 0.712 0.701
case10480_goc__api 99926 29816 196673 276 18 18 31 0.806 0.666 0.75
case13659_pegase 120495 38218 230046 0 122 107 81 2.21 3.29 2.69
case13659_pegase__api 120495 38218 230046 0 15 15 59 0.876 0.648 1.82
case20758_epigrids 182928 56275 355508 1881 18 18 32 1.43 1.43 1.28
case20758_epigrids__api 182928 56275 355508 1881 18 18 31 1.4 1.41 1.25
case19402_goc 184959 55077 364846 249 24 24 41 1.65 1.86 1.65
case19402_goc__api 184959 55077 364846 249 25 25 38 1.72 2.03 1.59
case19402_goc__sad 184959 55077 364846 249 22 - 55 1.61 - 2.09
case24464_goc 210481 63871 408258 348 22 - 26 1.82 - 1.52
case24464_goc__api 210481 63871 408258 348 21 21 25 1.76 2.09 1.41
case24464_goc__sad 210481 63871 408258 348 24 - 65 1.82 - 2.74
case30000_goc 213698 68919 399262 372 38 38 47 1.99 2.83 1.45
case30000_goc__api 213698 68919 399262 372 30 30 39 1.91 2.54 1.34
case30000_goc__sad 213698 68919 399262 372 25 25 83 1.65 1.89 2.24
case78484_epigrids 685984 211266 1334253 0 29 29 117 8.31 16.8 68.9
case78484_epigrids__api 685984 211266 1334253 0 26 26 106 7.89 15.2 93.8
case78484_epigrids__sad 685984 211266 1334253 0 30 30 120 8.37 17.6 68.7

### v3

@@ 4.3 Second-order cone programming
We compare our GPU implementation with the CPU-based Clarabel solver and MOSEK solver. We also include results of the MOSEK solver with pre-solve for comparison, which is denoted as Mosek* in the plots. The maximum termination time is again set to 11h. We benchmark the second-order cone relaxations of 120 problems from the PGLib-OPF library, where the number of second-order cones exceed $2000$.
@@
Overall, the GPU solver is several times faster than Rust-based CPU solver. This speedup causes ClarabelGPU to achieve a lower failure rate than ClarabelRs as it could solve more problems in under $10^{4}$ seconds each. However, note that ClarabelGPU fails on two examples that ClarabelRs successfully solves. This highlights the different numerical performance of the linear system solvers between cuDSS and faer-rs.
@@
Shifted GM Full Acc. 1.0 12.12 4.59 53.65
Low Acc. 1.0 4.16 3.91 7.29
Failure Rate (%) Full Acc. 1.7 43.6 10.3 89.7
Low Acc. 0.0 5.1 1.7 0.9
@@
We also consider the following multistage portfolio optimization problem:
minimize $\sum_{t=1}^{T}\left(-\mu_{t}^{\top}x_{t}+c\cdot\mathbf{1}^{\top}z_{t}+\gamma\cdot r_{t}\right)$ (21)
subject to $z_{1}\geq x_{1}-x_{0},\quad z_{1}\geq x_{0}-x_{1},$
$z_{t}\geq x_{t}-x_{t-1},\quad z_{t}\geq x_{t-1}-x_{t},\quad\forall t=2,\ldots,T,$
$\mathbf{1}^{\top}x_{1}=d+\mathbf{1}^{\top}x_{0},\quad\mathbf{1}^{\top}x_{t}=\mathbf{1}^{\top}x_{t-1},\quad\forall t=2,\ldots,T,$
$y_{t}=F_{t}x_{t},\quad\forall t=1,\ldots,T,$
$\left[\begin{array}[]{c}r_{t}\\ Uy_{t}\\ D_{\text{sqrt}}\odot x_{t}\end{array}\right]\in\mathcal{K}_{q}^{n+k+1},\quad\forall t=1,\ldots,T,$
with respect to $x$, $y$, $z$, and $r$. This problem is formulated over $T$ time periods, allocating wealth across $n$ assets while managing exposure to $k$ underlying factors. Here, $x_{t}$ represents asset allocations, $y_{t}$ represents factor exposures, $z_{t}$ represents trade volumes, and $r_{t}$ is a risk proxy. We impose additional constraints $x_{t}\in[0,0.1]^{n}$, $y_{t}\in[0,0.1]^{k}$, $z_{t}\in\mathbb{R}^{n}$, and $r_{t}\geq 0$. Our objective is a minimization over the trade-off between negative expected returns, linear transaction costs, and a second-order cone-based risk penalty scaled by risk aversion $\gamma$. The transaction constraints ensure that trade volumes ${z}_{t}$ correctly reflect changes in portfolio weights. Budget constraints enforce capital conservation: initial capital changes by a known inflow $d$, and total capital remains constant thereafter. Factor exposure is modeled via linear mappings ${y}_{t}=F_{t}{x}_{t}$. Risk $r_{t}$ is quantified through a second-order cone (SOC) constraint that captures both systematic risk $Uy_{t}$ and idiosyncratic risk $D_{\text{sqrt}}\odot x_{t}$ across each time period.
We benchmark multistage portfolio optimization problems with $n=5000,k=50$ and varying horizon $T$ in Table LABEL:table:parametric_programming_large_socp; ClarabelGPU achieved up to a 33x speedup over MOSEK and 4040x over ClarabelRs.
Table 2: Multi-stage portfolio optimization with varying horizons
iterations total time (s)
Horizon ClarabelGPU Mosek* ClarabelRs Mosek ClarabelGPU Mosek* ClarabelRs Mosek
5 21 15 21 14 0.882 1.24 12.1 1.65
10 22 15 22 15 1.79 3.87 53.7 4.63
15 22 14 22 12 2.8 7.49 99.5 7.42
20 22 15 22 11 3.91 12.5 167 13.2
25 24 18 24 14 9.82 20.2 245 19.8
30 23 12 23 13 11.3 22.1 312 28
@@ 4.4 Exponential cone programming
For exponential programming, we benchmark the entropy maximization problems [13] with varying dimensionality. The entropy maximization problem aims to maximize entropy over a probability distribution given a set of $m$ linear inequality constraints, which can be interpreted as bounds on the expectations of arbitrary functions. The problem is formulated as
@@
Each element of $A$ is generated from the distribution $A_{ij}\sim\mathcal{N}(0,n)$. Then, we set $b=Av/1^{T}v$ where $v\in\textbf{R}^{n}$ is generated randomly from $v_{i}\sim U[0,1]$, which ensures the problem is always feasible. We set $m$ to the nearest integer of $0.5n$ and vary value of $n$ from $2000$ to $10000$.
Table 3: Benchmarking for exponential cone programming
iterations total time (s)
Problem ClarabelGPU Mosek* ClarabelRs Mosek ClarabelGPU Mosek* ClarabelRs Mosek
entropy_max_2000 20 15 20 10 0.476 0.794 2.84 0.609
entropy_max_4000 20 16 20 9 1.4 3.02 14 2.28
entropy_max_6000 21 16 21 9 3.16 8.85 90.5 5.89
entropy_max_8000 21 16 21 10 5.98 17.7 104 12.1
entropy_max_10000 21 16 23 10 19.2 31.6 239 36.5
The benchmark results for these two problems with varying dimensionality are shown in Table LABEL:table:exponential_cone. Though the GPU acceleration of Clarabel is offset by nearly doubled number of iterations compared to MOSEK, we can still achieve more than 2 times of acceleration for the overall time. That is to say we can possibly achieve more acceleration in ClarabelGPU if we can improve the numerical stability to the same level of MOSEK on exponential cone programs. We find ClarabelGPU can benefit from GPU computation up to 10x times faster compared to ClarabelRs.
4.5 Semidefinite programming
Our results in Table LABEL:table:FEM_SDP show that ClarabelGPU achieves a speedup of $1.5\times$ to $4\times$ over MOSEK, both with and without presolve, and up to $10\times$ speedup compared to ClarabelRs. ClarabelGPU can only solve FELA_SDP_9263 to $1e^{-5}$ precision and is less numerical stable than Mosek* under the default setting, but increasing the static regularization $\delta_{s}$ to $1e^{-7}$ can fix the numerical issue.
Table 4: FEM with $3\times 3$-dimensional SDP constraints
iterations total time (s)
Problem ClarabelGPU Mosek* ClarabelRs Mosek ClarabelGPU Mosek* ClarabelRs Mosek
FELA_SDP_1048 23 24 - 22 2.41 6.75 - 2.84
OBEFM_SDP_1048 19 13 19 15 2.48 5.98 30.6 4.04
FELA_SDP_4444 30 31 - 38 11 40.8 - 19.4
OBEFM_SDP_4444 20 13 20 - 11.5 35.3 144 -
FELA_SDP_9263 - 33 31 34 - 55.1 167 41.5
OBEFM_SDP_9263 21 14 21 - 25.4 38.1 323 -
4.6 Parametric programming
Clarabel supports updating the coefficients $P,A,q,b$ without reinitializing the solver object. This feature is particularly advantageous for parametric programming, where the solver setup is required only once, and the symbolic factorization structure can be reused efficiently in subsequent solves [64]. Such capability is especially useful in applications like model predictive control and portfolio optimization.
For the multistage portfolio optimization problem in (21), we also split the total time into two parts, the setup time and the solve time, and report the ratio of them respectively in Table LABEL:table:time_ratio. Notably, the setup time for ClarabelGPU constitutes a larger proportion of the total time compared to its CPU counterpart. This suggests that, when the setup time dominates the total computational time, higher acceleration ratios can be achieved on a GPU when we need to solve multistage portfolio optimization multiple times with different parameters.
Table 5: Time ratio between setup and solve time on multistage portfolio optimization
ClarabelGPU Clarabel
Horizon setup time solve time setup time solve time
5 0.648 0.352 0.418 0.582
10 0.654 0.346 0.36 0.64
15 0.662 0.338 0.321 0.679
20 0.651 0.349 0.284 0.716
25 0.475 0.525 0.242 0.758
30 0.566 0.434 0.223 0.777
@@ 4.7 Mixed precision
We test the mixed precision setting on QPs, including portfolio optimization and Huber fitting problems from §4.2, for high accuracy level $\epsilon_{\text{feas}}=1e^{-8}$. We compare it with the standard GPU implementation of Float64 data type. The results in Table LABEL:table:mixed_large_qp demonstrate that employing a mixed-precision strategy can reduce solve time by up to a factor of 2 when the matrix $A$ becomes dense. However, we note that mixed precision is less numerically stable when solving an ill-conditioned KKT system (15). This can lead to an increased number of iterative refinement steps and longer solve times compared to full precision, or even failure to converge to the desired tolerance.
@@
iterations solve time (s)
@@
portfolio_optimization_n_5000 17 18 0.18 0.29
portfolio_optimization_n_10000 19 19 0.70 0.57
portfolio_optimization_n_15000 19 19 1.8 1.31
portfolio_optimization_n_20000 19 23 3.74 3.26
portfolio_optimization_n_25000 19 21 6.79 5.58
huber_fitting_n_5000 9 9 3.6 1.85
huber_fitting_n_10000 10 10 36.9 19.9
huber_fitting_n_15000 10 10 128 70.5
huber_fitting_n_20000 9 9 279 156
huber_fitting_n_25000 10 10 598 334
@@ 5 Conclusion
We have developed a GPU interior point solver for conic optimization. [^3] https://github.com/oxfordcontrol/Clarabel.jl/tree/CuClarabel In our implementation, we propose a mixed parallel computing strategy to process linear constraints with second-order cone, exponential cone, power cone and semidefinite cone constraints. Our GPU solver shows several times acceleration compared to state-of-the-art CPU conic solvers on many problems to high precision, such as QPs, SOCPs, exponential cone programs and SDPs.
Future research directions include extending support to general SDPs with PSD cones of varying dimensionalities. The proposed mixed parallel computing strategy for GPU implementation is also applicable to conic solvers based on first-order operator-splitting methods. This approach could improve GPU utilization, especially when sufficient computational resources are available to handle different cone classes in parallel. Additional performance gains may be achieved through kernel fusion, the use of CUDA graphs to reduce kernel launch overhead and overlapping more independent computation within interior-point methods.
@@ Acknowledgments
Yuwen Chen was supported by the EPSRC Impact Acceleration Account Fund and the National Centre of Competence in Research (NCCR) Automation, funded by the Swiss National Science Foundation (SNSF). Parth Nobel was supported in part by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-1656518. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The work of Stephen Boyd was supported by ACCESS (AI Chip Center for Emerging Smart Systems), sponsored by InnoHK funding, Hong Kong SAR.
Appendix A Scaling matrices
@@
The matrix $H(w)$ can be factorized as $H^{-1}(w)=W^{T}W$, and we set $H=H^{-1}(w)$ in (14). The factors $w,W$ are then computed following [52].
@@
Appendix B Parallel reduction for norm-$2$ computation
Algorithm 4 Parallel Reduction to Compute $\|u\|^{2}$
[…]
0: Scalar $\texttt{norm2}=\sum_{j=1}^{n}u_{j}^{2}$
1:
2:inline function _reduction_norm2(u, n)
3:   Allocate shared_mem[1 .. T] = 0 // T = threads per block
4:   $j\leftarrow$ threadIdx.x
5:   while $j\leq n$ then
6:     shared_mem[j] $\leftarrow\texttt{shared\_mem[j]}+u[j]^{2}$ // Julia is 1-indexed
7:   end
8:   _syncthreads()
@@ Appendix B Parallel reduction for norm-$2$ computation
10: // Set $s$ to the largest number of power 22 that is smaller than $n$
11:   s = prevpow(2,n1)(2,n-1)
12:   while $s>1$
13:     if $(j\leq s\ \&\&\ j+s\leq n)$ then
14:       shared_mem[j] $\leftarrow$ shared_mem[j] ++ shared_mem[j + s]
15:     end if
16:     _syncthreads()
17:     $s>>1$   //Divide $s$ by 22
18:   end for
19:   if $j==1$ then
20:     norm2 $\leftarrow$ shared_mem[1]
21:   end if
22:end
Appendix C Detailed benchmark results